# Interpreting Graph Inference with Skyline Explanations

### v3

@@ VI Experimental Study
We conduct experiments to evaluate the effectiveness, efficiency, and scalability of our solutions. Our algorithms are implemented in Python 3.10.14 by PyTorch-Geometric framework. All experiments are conducted on a Linux system equipped with AMD Ryzen 9 5950X CPU, an NVIDIA GeForce RTX 3090, and 32 GB of RAM. Our code and data are made available at [61].
@@ VI-A Experimental Setup
Datasets. Table III summarizes the datasets. (1) $\mathsf{Cora}$ [56] and $\mathsf{PubMed}$ [57] are citation networks where nodes (papers) are classified by category using keyword-based features. (2) $\mathsf{FacebookPage}$ ($\mathsf{FB}$[58] contains verified Facebook pages (nodes) linked by mutual likes, with features from page descriptions; the task is multi-class page classification. (3) $\mathsf{AmazonComputer}$ ($\mathsf{AMZ}$[59] is a co-purchase network of computer products with review-based features; the task is product category classification. (4) $\mathsf{OGBN\_arxiv}$ ($\mathsf{arxiv}$[60] is a citation network of CS papers with 128-d word-embedding features, classified by subject area. (5) $\mathsf{BA\_Shapes}$ ($\mathsf{BA}$[6] is a synthetic BA graph with house motifs; house nodes serve as human-crafted ground-truth explanations. (6) $\mathsf{Tree\_Cycles}$ ($\mathsf{Tree}$[6] is a synthetic tree graph with cycle motifs; cycle nodes serve as human-crafted ground-truth explanation [5].
GNN Classifiers. We employ three classes of mainstream $\mathsf{GNNs}$: (1) Graph convolutional network ($\mathsf{GCN}$[29], one of the classic message-passing $\mathsf{GNNs}$; (2) Graph attention networks ($\mathsf{GAT}$[30] leverage attention mechanisms to dynamically weigh the importance of a node’s neighbors during inference; and (3) Graph isomorphism networks ($\mathsf{GIN}$[31] with enhanced expressive power up to the Weisfeiler-Lehman graph isomorphism tests.
@@
We compare against several state-of-the-art $\mathsf{GNN}$ explainers. (1) GNNExplainer ($\mathsf{GExp}$) learns edge and feature masks by maximizing mutual information between predictions on the original and masked graphs [6]. (2) PGExplainer ($\mathsf{PGExp}$) trains an MLP as a mask generator on node embeddings, also optimizing mutual information [10]. (3) $\mathsf{CF\textsuperscript{2}}$ learns feature and edge masks by optimizing a weighted combination of factual and counterfactual measures [8]. (4) $\mathsf{MOExp}$ produces Pareto-optimal explanations balancing simulatability and counterfactual relevance [11]; without constraint $k$, it may return up to $\sim$200 subgraphs. (5) SAME develops structure-aware Shapley-based explanations via connected substructures [34].
@@
Evaluation Metrics. For all datasets and $\mathsf{GNNs}$, we select three common explainability measures ($\Phi$): $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$. In addition, we adopt the explanation accuracy [6] ($\mathsf{acc}$) for the synthetic $\mathsf{Tree\_Cycles}$ dataset. $\mathsf{acc}$ measures the proportion of human-crafted ground-truth nodes present in the explanatory subgraph. For computational simplicity, we transformed each minimization measure ($\mathsf{conc}$}, $\mathsf{fdl}$$^{-}$) into its inverse counterpart, so that all measures are expressed as maximization objectives. As most $\mathsf{GNN}$ explainers are not designed to generate explanations for multiple explainability measures, for a fair comparison, we employ three quality indicators (QIs)  [62, 63, 64, 65]. These quality indicators are widely-used to measure how good each result is in a multi-objective manner. Consider a set of $n$ $\mathsf{GNN}$ explainers, where each explainer $A$ reports a set of explanatory subgraphs ${\mathcal{G}}_{A}$.
(1) QI-1: Integrated Preference Function (IPF) (Higher is better) [66]. IPF score unifies and compares the quality of non-dominated set solutions with a weighted linear sum function. We define a normalized IPF of an explanation ${\mathcal{G}}_{A}$ from each $\mathsf{GNN}$ explainer $A$ with a normalized single-objective score:
@@
(2) QI-2: Inverted Generational Distance (IGD) (Lower is better)  [67, 62], a most commonly used distance-based QI. It measures the distance from each solution to a reference set that contains top data points with theoretically achievable “ideal” values. We introduce IGD for explanations as follows. (a) We define a universal space $\mathcal{G}$} = $\bigcup_{i\in[1,n]}{\mathcal{G}}_{i}$ from all the participating $\mathsf{GNN}$ explainers, and for each $\phi\in\Phi$, induces a reference set ${\mathcal{G}}_{\phi}^{k}\in{\mathcal{G}}$ with explanatory subgraphs having the top-$k$ values in $\phi$. (b) The normalized IGD of an explanation ${\mathcal{G}}_{A}$ from an explainer $A$ is defined as:
@@
$d(\cdot)$ is Euclidean distance function following [62, 63].
(3) QI-3: Maximum Spread (MS) (Higher is better) [62]. $\mathsf{MS}$ is a widely-adopted spread indicator that quantifies the range of the minimum and maximum values a solution can achieve in each objective. For a fair comparison, we introduce a normalized $\mathsf{MS}$ score using reference sets in QI-2. For each measure $\phi\in\Phi$, and an explanation ${\mathcal{G}}_{A}$, its normalized MS score on $\phi$ is computed as:
@@ VI-B Experimental Results
QI-1: IPF Scores. We report IPF scores (bigger is better) for all $\mathsf{GNN}$ explainers in Figure 3(a) with $\mathsf{GCNs}$ and kk=$10$. Additional explainability results with varying $k$ are given in the full version [47]. (1) In general, $\mathsf{ASX}$-$\mathsf{OP}$ outperforms a majority of competitors in multi-objective explainability evaluation metrics. For example, on $\mathsf{Cora}$, $\mathsf{ASX}$-$\mathsf{OP}$ outperforms $\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, and $\mathsf{SAME}$ in IPF scores by 1.34, 1.51, 2.81, 1.79, and 1.23 times, respectively. $\mathsf{DSX}$ and $\mathsf{ASX}$-$\mathsf{I}$ achieve comparable performance with $\mathsf{ASX}$-$\mathsf{OP}$. (2) We also observe that $\mathsf{GExp}$ and $\mathsf{PGExp}$ are sensitive as the datasets vary. For example, on $\mathsf{FacebookPage}$, both show a significant change in IPF scores. In contrast, $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ consistently achieve top IPF scores over all datasets. This verifies their robustness in generating high-quality explanations over different data sources.
@@
Exp-3: Scalability. We report the impact of critical factors (i.e., number of explanatory subgraphs $k$, $\mathsf{GNN}$ classes, and number of $\mathsf{GNN}$ layers $L$) on the scalability of skyline explanation generation, using $\mathsf{Cora}$ dataset. Additional results about effectiveness by varying factors are in full version [47].
@@
In Figure 5(b), we visualize the distribution of explanatory subgraphs for $v_{1}$ in Figure 5(a). Each point represents a subgraph with 3D coordinates given by $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$} scores. The gray dots denote the verified explanatory subgraph in the interpretable space, while baseline explainers ($\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, $\mathsf{SAME}$) are shown as diamonds. Our skyline explanations appear as red crosses, with their convex hull illustrating dominance coverage. We observe that skyline explanations span most of the interpretable space, offering comprehensive and diverse explanations, whereas baseline explanations cluster in smaller regions, missing diversity [18].
@@ VII Conclusion
We proposed a class of skyline explanations that optimize multiple explainability measures for interpreting GNN predictions. We showed that generating skyline explanations, viewed as an explanatory query processing problem, is nontrivial even with polynomially bounded input space and three measures. To address this, we developed sequential, parallel, and diversified algorithms with provable Pareto-optimality guarantees. Extensive experiments demonstrate that our methods are practical for large-scale graphs and GNN-based classification, producing more comprehensive explanations than state-of-the-art GNN explainers.
@@ VII-A Details of Algorithms and Procedures
Next, we identify the skyline $\overline{s}$ with the smallest ${\mathcal{D}}(\overline{s})$, replace $\overline{s}$ with $s$, only when such replacement makes the $\mathcal{D}$} increased by a factor of $\frac{1}{k}$. If swapping is conducted, the new skyline, i.e. $s$, connects directed edges to the dominating states based on bitvector $B(s)$, and we remove the swapped skyline and its corresponding edges. The above swap strategy ensures a $\frac{1}{4}$-approximation ratio [49].
@@
Approximability and Time cost of $\mathsf{DSX}$. We can verify that ${\sf NCS}(\cdot)$ and ${\sf CD}(\cdot)$ are submodular functions. Consider procedure updateDSX upon the arrival, at any time, of a new verified candidate $G_{s}$. Given that $|{\mathcal{G}}_{\epsilon}|\leq k$ is a hard constraint, we reduce the diversified evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the Streaming Submodular Maximization problem [53]. The problem maintains a size-$k$ set that optimizes a submodular function over a stream of data objects. The diversification function $\sf DivS$} is submodular, as it combines two submodular components: $\sf NCS$} and the summation over $\sf CD$}. Specifically, $\sf NCS$}, representing node coverage, is a submodular function bounded within $[0,N]$, with each incremental gain diminishing as more nodes are covered. In parallel, $\sf CD$} denotes the cosine distance between two embeddings, which is bounded in $[0,1]$. Since pairwise cosine distance is a non-negative and bounded measure, the summation over $\sf CD$} across a set remains submodular. Consequently, their combination in $\sf DivS$} preserves submodularity. $\mathsf{DSX}$ adopts a greedy increment policy by including a candidate in ${\mathcal{G}}_{\epsilon}$ with the new candidate $G_{s}$ only when this leads to a marginal gain greater than $\frac{(1+\epsilon)/2-f(S)}{k-|S|}$, where $S$ is the current set and $f(\cdot)$ is a submodular function. This is consistent with an increment policy that ensures a ($\frac{1}{2}-\epsilon$)-approximation in [53]. Since $\mathsf{DSX}$ follows the same process as $\mathsf{ASX}$-$\mathsf{OP}$ but only differs in replacement policy with same time cost, the cost of $\mathsf{DSX}$ is also $O(|\zeta^{\prime}|(\log\frac{r_{\Phi}}{\epsilon})^{|\Phi|}+|\zeta^{\prime}|L|G^{L}(v_{t})|)$.
@@
(2) Influential Node Prioritization. Firstly, since edge information is computed with respect to individual test nodes, any shared edge is attributed to the first node that processes it. To ensure this shared information is representative of other test nodes within the same thread, it is important to prioritize nodes that are more influential or centrally located among the assigned subset. To this end, we propose an influential node prioritization strategy that ranks test nodes within each thread, ensuring that nodes with higher centrality are processed earlier. Each thread begins by applying the $k$-core decomposition algorithm [68] to assign a core number to each node, reflecting the highest $j$ such that the node belongs to the $j$-core. Test nodes are then ranked in descending order of core number, prioritizing more “central” nodes in the thread. This ordering ensures that edge information from structurally important nodes is shared early, maximizing reuse.
@@
Example 6: As illustrated in Figure 8, six $\mathsf{SXQ}$ queries have been issued with specified output nodes $v_{1}$ to $v_{6}$. Following the intra-similarity analysis (procedure $\mathsf{CP}$), the query workload is partitioned into three clusters $\{v_{1},v_{3}\}$, $\{v_{4},v_{5}\}$ and $\{v_{2},v_{6}\}$ (the latter two are not shown), by maximizing the intra-cluster Jaccard similarity of their $2$-hop neighbors. The nodes in the same cluster are sorted by their core numbers by adapting $k$-core decomposition algorithm [68]. In this case, $v_{1}$ is scheduled to be processed first, followed by $v_{3}$. During the processing, all threads consistently invoke $\mathsf{ASX}$-$\mathsf{OP}$, yet share the edge information asynchronously, via the shared edge information table. For example, once the edges in the neighbors of $v_{1}$ are verified, its auxiliary information is shared for fast local computation without recomputation. For instance, $e_{17}$ is skipped for $v_{3}$ since it has $e_{17}$ in its neighbor graphs. $\Box$

### v4

@@ VI Experimental Study
We conduct experiments to evaluate the effectiveness, efficiency, and scalability of our solutions. Our algorithms are implemented in Python 3.10.14 by PyTorch-Geometric framework. All experiments are conducted on a Linux system equipped with AMD Ryzen 9 5950X CPU, an NVIDIA GeForce RTX 3090, and 32 GB of RAM. Our code and data are made available at [code_base].
@@ VI-A Experimental Setup
Datasets. Table III summarizes the datasets. (1) $\mathsf{Cora}$ [cora] and $\mathsf{PubMed}$ [pubmed] are citation networks where nodes (papers) are classified by category using keyword-based features. (2) $\mathsf{FacebookPage}$ ($\mathsf{FB}$[facebook] contains verified Facebook pages (nodes) linked by mutual likes, with features from page descriptions; the task is multi-class page classification. (3) $\mathsf{AmazonComputer}$ ($\mathsf{AMZ}$[shchur2018pitfalls] is a co-purchase network of computer products with review-based features; the task is product category classification. (4) $\mathsf{OGBN\_arxiv}$ ($\mathsf{arxiv}$[hu2020open] is a citation network of CS papers with 128-d word-embedding features, classified by subject area. (5) $\mathsf{BA\_Shapes}$ ($\mathsf{BA}$[ying2019gnnexplainer] is a synthetic BA graph with house motifs; house nodes serve as human-crafted ground-truth explanations. (6) $\mathsf{Tree\_Cycles}$ ($\mathsf{Tree}$[ying2019gnnexplainer] is a synthetic tree graph with cycle motifs; cycle nodes serve as human-crafted ground-truth explanation [yuan2022explainability].
GNN Classifiers. We employ three classes of mainstream $\mathsf{GNNs}$: (1) Graph convolutional network ($\mathsf{GCN}$[kipf2016semi], one of the classic message-passing $\mathsf{GNNs}$; (2) Graph attention networks ($\mathsf{GAT}$[gat] leverage attention mechanisms to dynamically weigh the importance of a node’s neighbors during inference; and (3) Graph isomorphism networks ($\mathsf{GIN}$[gin] with enhanced expressive power up to the Weisfeiler-Lehman graph isomorphism tests.
@@
We compare against several state-of-the-art $\mathsf{GNN}$ explainers. (1) GNNExplainer ($\mathsf{GExp}$) learns edge and feature masks by maximizing mutual information between predictions on the original and masked graphs [ying2019gnnexplainer]. (2) PGExplainer ($\mathsf{PGExp}$) trains an MLP as a mask generator on node embeddings, also optimizing mutual information [luo2020parameterized]. (3) $\mathsf{CF\textsuperscript{2}}$ learns feature and edge masks by optimizing a weighted combination of factual and counterfactual measures [tan2022learning]. (4) $\mathsf{MOExp}$ produces Pareto-optimal explanations balancing simulatability and counterfactual relevance [liu2021multi]; without constraint $k$, it may return up to $\sim$200 subgraphs. (5) SAME develops structure-aware Shapley-based explanations via connected substructures [ye2023same].
@@
Evaluation Metrics. For all datasets and $\mathsf{GNNs}$, we select three common explainability measures ($\Phi$): $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$. In addition, we adopt the explanation accuracy [ying2019gnnexplainer] ($\mathsf{acc}$) for the synthetic $\mathsf{Tree\_Cycles}$ dataset. $\mathsf{acc}$ measures the proportion of human-crafted ground-truth nodes present in the explanatory subgraph. For computational simplicity, we transformed each minimization measure ($\mathsf{conc}$}, $\mathsf{fdl}$$^{-}$) into its inverse counterpart, so that all measures are expressed as maximization objectives. As most $\mathsf{GNN}$ explainers are not designed to generate explanations for multiple explainability measures, for a fair comparison, we employ three quality indicators (QIs)  [li2019quality, qi2, qi3, qi4]. These quality indicators are widely-used to measure how good each result is in a multi-objective manner. Consider a set of $n$ $\mathsf{GNN}$ explainers, where each explainer $A$ reports a set of explanatory subgraphs ${\mathcal{G}}_{A}$.
(1) QI-1: Integrated Preference Function (IPF) (Higher is better) [carlyle2003quantitative]. IPF score unifies and compares the quality of non-dominated set solutions with a weighted linear sum function. We define a normalized IPF of an explanation ${\mathcal{G}}_{A}$ from each $\mathsf{GNN}$ explainer $A$ with a normalized single-objective score:
@@
(2) QI-2: Inverted Generational Distance (IGD) (Lower is better)  [coello2004study, li2019quality], a most commonly used distance-based QI. It measures the distance from each solution to a reference set that contains top data points with theoretically achievable “ideal” values. We introduce IGD for explanations as follows. (a) We define a universal space $\mathcal{G}$} = $\bigcup_{i\in[1,n]}{\mathcal{G}}_{i}$ from all the participating $\mathsf{GNN}$ explainers, and for each $\phi\in\Phi$, induces a reference set ${\mathcal{G}}_{\phi}^{k}\in{\mathcal{G}}$ with explanatory subgraphs having the top-$k$ values in $\phi$. (b) The normalized IGD of an explanation ${\mathcal{G}}_{A}$ from an explainer $A$ is defined as:
@@
$d(\cdot)$ is Euclidean distance function following [li2019quality, qi2].
(3) QI-3: Maximum Spread (MS) (Higher is better) [li2019quality]. $\mathsf{MS}$ is a widely-adopted spread indicator that quantifies the range of the minimum and maximum values a solution can achieve in each objective. For a fair comparison, we introduce a normalized $\mathsf{MS}$ score using reference sets in QI-2. For each measure $\phi\in\Phi$, and an explanation ${\mathcal{G}}_{A}$, its normalized MS score on $\phi$ is computed as:
@@ VI-B Experimental Results
QI-1: IPF Scores. We report IPF scores (bigger is better) for all $\mathsf{GNN}$ explainers in Figure 3(a) with $\mathsf{GCNs}$ and kk=$10$. Additional explainability results with varying $k$ are given in Appendix. (1) In general, $\mathsf{ASX}$-$\mathsf{OP}$ outperforms a majority of competitors in multi-objective explainability evaluation metrics. For example, on $\mathsf{Cora}$, $\mathsf{ASX}$-$\mathsf{OP}$ outperforms $\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, and $\mathsf{SAME}$ in IPF scores by 1.34, 1.51, 2.81, 1.79, and 1.23 times, respectively. $\mathsf{DSX}$ and $\mathsf{ASX}$-$\mathsf{I}$ achieve comparable performance with $\mathsf{ASX}$-$\mathsf{OP}$. (2) We also observe that $\mathsf{GExp}$ and $\mathsf{PGExp}$ are sensitive as the datasets vary. For example, on $\mathsf{FacebookPage}$, both show a significant change in IPF scores. In contrast, $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ consistently achieve top IPF scores over all datasets. This verifies their robustness in generating high-quality explanations over different data sources.
@@
Exp-3: Scalability. We report the impact of critical factors (i.e., number of explanatory subgraphs $k$, $\mathsf{GNN}$ classes, and number of $\mathsf{GNN}$ layers $L$) on the scalability of skyline explanation generation, using $\mathsf{Cora}$ dataset. Additional results about effectiveness by varying factors are in Appendix.
@@
In Figure 5(b), we visualize the distribution of explanatory subgraphs for $v_{1}$ in Figure 5(a). Each point represents a subgraph with 3D coordinates given by $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$} scores. The gray dots denote the verified explanatory subgraph in the interpretable space, while baseline explainers ($\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, $\mathsf{SAME}$) are shown as diamonds. Our skyline explanations appear as red crosses, with their convex hull illustrating dominance coverage. We observe that skyline explanations span most of the interpretable space, offering comprehensive and diverse explanations, whereas baseline explanations cluster in smaller regions, missing diversity [das1997closer].
@@ VII Conclusion
We proposed a class of skyline explanations that optimize multiple explainability measures for interpreting GNN predictions. We showed the hardness of explanatory query processing problem. To address this, we developed sequential, parallel, and diversified algorithms with provable Pareto-optimality guarantees. Extensive experiments demonstrate that our methods are practical for large-scale graphs and GNN-based classification, producing more comprehensive explanations than state-of-the-art GNN explainers.
Acknowledgment. Dazhuo Qiu and Arijit Khan acknowledge support from the Novo Nordisk Foundation grant NNF22OC0072415. Haolai Che and Yinghui Wu acknowledge support from the NSF OAC-2104007.
(a) IPF varying $k$
(b) IGD varying $k$
(c) IPF varying $\mathsf{GNNs}$
(d) IGD varying $L$
Figure 7: Impact of factors on the effectiveness of explanations
@@ VII-A Details of Algorithms and Procedures
Next, we identify the skyline $\overline{s}$ with the smallest ${\mathcal{D}}(\overline{s})$, replace $\overline{s}$ with $s$, only when such replacement makes the $\mathcal{D}$} increased by a factor of $\frac{1}{k}$. If swapping is conducted, the new skyline, i.e. $s$, connects directed edges to the dominating states based on bitvector $B(s)$, and we remove the swapped skyline and its corresponding edges. The above swap strategy ensures a $\frac{1}{4}$-approximation ratio [onlinemkc].
@@
Approximability and Time cost of $\mathsf{DSX}$. We can verify that ${\sf NCS}(\cdot)$ and ${\sf CD}(\cdot)$ are submodular functions. Consider procedure updateDSX upon the arrival, at any time, of a new verified candidate $G_{s}$. Given that $|{\mathcal{G}}_{\epsilon}|\leq k$ is a hard constraint, we reduce the diversified evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the Streaming Submodular Maximization problem [ssm]. The problem maintains a size-$k$ set that optimizes a submodular function over a stream of data objects. The diversification function $\sf DivS$} is submodular, as it combines two submodular components: $\sf NCS$} and the summation over $\sf CD$}. Specifically, $\sf NCS$}, representing node coverage, is a submodular function bounded within $[0,N]$, with each incremental gain diminishing as more nodes are covered. In parallel, $\sf CD$} denotes the cosine distance between two embeddings, which is bounded in $[0,1]$. Since pairwise cosine distance is a non-negative and bounded measure, the summation over $\sf CD$} across a set remains submodular. Consequently, their combination in $\sf DivS$} preserves submodularity. $\mathsf{DSX}$ adopts a greedy increment policy by including a candidate in ${\mathcal{G}}_{\epsilon}$ with the new candidate $G_{s}$ only when this leads to a marginal gain greater than $\frac{(1+\epsilon)/2-f(S)}{k-|S|}$, where $S$ is the current set and $f(\cdot)$ is a submodular function. This is consistent with an increment policy that ensures a ($\frac{1}{2}-\epsilon$)-approximation in [ssm]. Since $\mathsf{DSX}$ follows the same process as $\mathsf{ASX}$-$\mathsf{OP}$ but only differs in replacement policy with same time cost, the cost of $\mathsf{DSX}$ is also $O(|\zeta^{\prime}|(\log\frac{r_{\Phi}}{\epsilon})^{|\Phi|}+|\zeta^{\prime}|L|G^{L}(v_{t})|)$.
@@
(2) Influential Node Prioritization. Firstly, since edge information is computed with respect to individual test nodes, any shared edge is attributed to the first node that processes it. To ensure this shared information is representative of other test nodes within the same thread, it is important to prioritize nodes that are more influential or centrally located among the assigned subset. To this end, we propose an influential node prioritization strategy that ranks test nodes within each thread, ensuring that nodes with higher centrality are processed earlier. Each thread begins by applying the $k$-core decomposition algorithm [core_decomp] to assign a core number to each node, reflecting the highest $j$ such that the node belongs to the $j$-core. Test nodes are then ranked in descending order of core number, prioritizing more “central” nodes in the thread. This ordering ensures that edge information from structurally important nodes is shared early, maximizing reuse.
@@
Example 6: As illustrated in Figure 8, six $\mathsf{SXQ}$ queries have been issued with specified output nodes $v_{1}$ to $v_{6}$. Following the intra-similarity analysis (procedure $\mathsf{CP}$), the query workload is partitioned into three clusters $\{v_{1},v_{3}\}$, $\{v_{4},v_{5}\}$ and $\{v_{2},v_{6}\}$ (the latter two are not shown), by maximizing the intra-cluster Jaccard similarity of their $2$-hop neighbors. The nodes in the same cluster are sorted by their core numbers by adapting $k$-core decomposition algorithm [core_decomp]. In this case, $v_{1}$ is scheduled to be processed first, followed by $v_{3}$. During the processing, all threads consistently invoke $\mathsf{ASX}$-$\mathsf{OP}$, yet share the edge information asynchronously, via the shared edge information table. For example, once the edges in the neighbors of $v_{1}$ are verified, its auxiliary information is shared for fast local computation without recomputation. For instance, $e_{17}$ is skipped for $v_{3}$ since it has $e_{17}$ in its neighbor graphs. $\Box$